OpenStudy (anonymous):

Indefinite integral wrt x of \[x^{2}\sqrt{0.5x+3} \]?

OpenStudy (anonymous):

what is wrt?

OpenStudy (anonymous):

with respect to

OpenStudy (anonymous):

i think you should take couple integration by parts: 1) u=x^2 du=2x dv=(.5x+3)^1/2; v=4/3 (2/3 x +3)^3/2 i probably one more time to reduce the power of x^2

OpenStudy (anonymous):

first integration by parts would give: \[4/3 x^2 (1/2 x +3)^{3/2} -8/3 \int\limits_{ }^{}x*(1/2 x+3)^{3/2}dx=\] I think one more time would do it... :)

OpenStudy (anonymous):

did you finished it ? need help?

OpenStudy (anonymous):

Put u = 0.5x+3 (x = 2u-6) , du = 0.5 -> Int 2(2u-6)^2 root u du Int(8u^(5/2) -48u^(3/2)+72u^(1/2) du and sub back for u

OpenStudy (anonymous):

I got final result as (but I would check coefficients... was rushing...): \[=4/3 *x ^{2}(1/2 x +3)^{3/2}-32/15 *x(1/2 x+3)^{5/2} +\] \[+128/105 (1/2 x +3)^{7/2} + const\]

OpenStudy (anonymous):

Basically right, as you say coefficients are off a bit.